OpenStudy (anonymous):

The ratio between the radius of the base and the height of the cylinder is 2:3.if its volume is 1617cm^3,the total surface area of the cylinder is?

Directrix (directrix):

Do know the formula for the volume of a right circular cylinder? If you do, I won't have to look for it. :)

Directrix (directrix):

I know how to do the problem. I just need that formula.

OpenStudy (anonymous):

pi r^2h

Directrix (directrix):

Ok. Let h = 3x and r = 2x. r/h = 2/3 so we're okay with that. V = pi( r^2)h 1617 = pi * (2x)(2x) ( 3x) Solve for x and post here.

Directrix (directrix):

1617 = pi * 12 x^3 Now, what's next?

OpenStudy (anonymous):

x=3.5

Directrix (directrix):

Based on your work, then h = 3*(3.5) and r = 2(3.5). So, what are the values of h and r?

OpenStudy (anonymous):

h=10.5 r=7

Directrix (directrix):

Total Area of cylinder = 2 pi r h + 2 B where B is the area of the circular base. TA = 2 pi r h + 2 pi r ^2 Fill in for r and h and crank out the answer. Post here.

OpenStudy (anonymous):

T.A=770

Directrix (directrix):

square cm. Yes.